OpenStudy (anonymous):

Find all solutions to the equation. (sin x)(cos x) = 0 a.{npi|n=0,1,2,...} b.{(pi/2)+npi,npi|n=0,1,2,...} c.{(pi/2)+2npi|n=0,1,2,...} d.{(pi/2)+npi|n=0,1,2,...}

OpenStudy (anonymous):

sinx cosx = 1/2sin2x =0 iff 2x = pi/2 + k pi or x = pi/4 +k/2 pi

OpenStudy (anonymous):

@Hoa so which of the following on the option is it? O.o

OpenStudy (anonymous):

@hannahg for all x satisfy the condition, those are the solution of the equation, i think

OpenStudy (anonymous):

@Hoa so the answer will be an a??